# Operator algebra of the information-disturbance tradeoff in quantum measurements

### v1

@@ I Introduction
It is well known that there is an information-disturbance tradeoff in quantum mechanics. Any information gain due to a measurement of the system also implies a disturbance of that system [1, 2]. In terms of operators, a quantum measurement necessarily changes states that are not eigenstates of the measurement operator. However, the precise evaluation of this tradeoff using concepts of information theory is surprisingly difficult [3, 4, 5, 6, 7]. Existing literature employs tools from quantum information theory such as mutual information, quantum discord, and mean estimation fidelity to bound the tradeoff between information gain and the disturbance introduced to the system during measurement [9, 10, 11]. These tools tend to focus on the information content of quantum states, making it somewhat difficult to identify the role of the measurement process itself in the definition of the tradeoff. Here, we propose to shift the focus to the operators that describe the measurement process, introducing an analysis of the operator algebra that defines the information-disturbance tradeoff.
@@ II Operator expressions of information and disturbance
A minimally disturbing measurement process that distinguishes a set of orthogonal states $\{\left|a\right\rangle\}$ can be represented by a self-adjoint measurement operator $\hat{M}_{m}$, where $m$ represents the outcome of the measurement related to $a$ by the conditional probabilities $p(m|a)$. The spectral decomposition of the measurement operator is given by
@@
The projection operators $\left|a\right\rangle\thinspace\left\langle a\right|$ represent the causal relation between the target observable described by the eigenstates $\left|a\right\rangle$ and the measurement outcome $m$. In this sense, the operator representation is biased in favor of the information extraction represented by the eigenvalues $\sqrt{p(m|a)}$. At the same time, the operator necessarily changes the quantum state that it acts upon by modifying the coherences between the eigenstate components $\left|a\right\rangle$. This modification of the coherences between eigenstates describes the necessary disturbance of the state caused by the extraction of information described by the conditional probabilities $p(m|a)$. It follows that we can identify the physics of this disturbance by changing the representation of the operators $\hat{M}_{m}$ to a description that identifies the precise change of the coherences between the eigenstates $\left|a\right\rangle$.
@@ V Conclusions
The information gain described by minimally disturbing operators $\hat{M}_{m}$ is described by conditional probabilities $p(m|a)$ that completely characterize the operator. It is therefore possible to characterize the information-disturbance trade-off by transforming the spectral decomposition of the operator $\hat{M}_{m}$ into a superposition of unitary operations that represent the unavoidable disturbance. As we have shown above, the results of this transformation provide deeper insights into the relation between information and disturbance as an intrinsic feature of the measurement process. Eq. (14) describes the experimentally observable bound on information gained about $a$ for a specific pattern of disturbance observed in the complementary property $b$. When characterizing a quantum channel, the observed disturbance of the inputs $\{\left|b\right\rangle\}$ can be used to put an upper limit on the possible leakage of information encoded in $\{\left|a\right\rangle\}$. This limit may be useful in a quick assessment of quantum cryptography protocols, providing the likelihoods of identifying eavesdroppers in BB84 type protocols where the information is encoded in two complementary basis systems $\{\left|a\right\rangle\}$ and $\{\left|b\right\rangle\}$ [18].
It should be emphasized that the complementary basis $\{\left|b\right\rangle\}$ appears in the analysis as a description of the distinguishable effects of unitaries $\hat{U}(k)$ with eigenstates $\left|a\right\rangle$. The relation between the basis sets $\{\left|a\right\rangle\}$ and $\{\left|b\right\rangle\}$ thus originates from the relation between phase changes and experimentally observable changes in quantum mechanics. It might be worth noting that the same fundamental relation between phases and observable output is expressed by the quantum Fourier transform that appears in Shor’s algorithm [19]. In general, the quantum formalism relates information and dynamics by describing quantum processes as Hilbert space operators, where the same algebra is used for the representation of information gain by projection operators and for the representation of disturbance by unitary operators. The fundamental relation between information and disturbance formulated in this work thus captures a fundamental aspect of quantum mechanics with no correspondence in either classical physics or classical information theory.

### v2

# An operator-based bound on information and disturbance in quantum measurements

@@ I Introduction
It is well known that there is an information-disturbance tradeoff in quantum mechanics. Any information gain due to a measurement of the system also implies a disturbance of that system [1, 2]. In terms of operators, a quantum measurement necessarily changes states that are not eigenstates of the measurement operator. However, the precise evaluation of this tradeoff using concepts of information theory is surprisingly difficult [3, 4, 5, 6, 7, 8]. Existing literature employs tools from quantum information theory such as mutual information, quantum discord, and mean estimation fidelity to bound the tradeoff between information gain and the disturbance introduced to the system during measurement [9, 10, 11]. These tools tend to focus on the information content of quantum states, making it somewhat difficult to identify the role of the measurement process itself in the definition of the tradeoff. Here, we propose to shift the focus to the operators that describe the measurement process, introducing an analysis of the operator algebra that defines the information-disturbance tradeoff.
@@ II Operator expressions of information and disturbance
We consider a measurement designed to extract information about an observable $\hat{A}$ with eigenstates $\{\left|a\right\rangle\}$ whilst avoiding any unnecessary dynamical disturbance of the system. Such measurements are naturally described as quantum non-demolition (QND) measurements, for which the back action does not induce transitions between different eigenstates of the measured observable. In this setting, the measurement operators are diagonal in the $\{\left|a\right\rangle\}$ basis and the associated POVM elements commute.
In general, measurement operators need not be self-adjoint, as they may include additional unitary back action. However, any Kraus operator $\hat{K}$ admits a decomposition $\hat{K}=\hat{U}\hat{M}$, where $\hat{M}=\sqrt{\hat{K}^{\dagger}\hat{K}}$ is positive and $\hat{U}$ is unitary. Here, the positive operator $\hat{M}$ uniquely determines the conditional probabilities of the measurement outcomes, whereas the unitary operator $\hat{U}$ represents an additional back action which does not affect these probabilities and therefore does not contribute to the information extracted by the measurement. As shown in the context of measurement-based feedback control, such unitary back action can in principle be eliminated, allowing the realization of minimally disturbing measurements characterized by self-adjoint operators only [18].
Motivated by these insights, we restrict attention to minimally disturbing measurements described by self-adjoint operators $\hat{M}_{m}$ with eigenstates $\{\left|a\right\rangle\}$ and real eigenvalues $\sqrt{p(m|a)}$, where $m$ represents the outcome of the measurement related to $a$ by the conditional probabilities $p(m|a)$. The spectral decomposition of the measurement operator is given by
@@
The projection operators $\left|a\right\rangle\thinspace\left\langle a\right|$ represent the causal relation between the target observable described by the eigenstates $\left|a\right\rangle$ and the measurement outcome $m$. In this sense, the operator representation is biased in favor of the information extraction represented by the eigenvalues $\sqrt{p(m|a)}$. At the same time, the operator necessarily changes the quantum state that it acts upon by modifying the coherences between the eigenstate components $\left|a\right\rangle$. This modification of the coherences between eigenstates describes the necessary disturbance of the state caused by the extraction of information described by the conditional probabilities $p(m|a)$. It follows that we can identify the physics of this disturbance by changing the representation of the operators $\hat{M}_{m}$ to a description that identifies the precise change of the coherences between the eigenstates $\left|a\right\rangle$. To make this disturbance explicit, it is useful to express $\hat{M}_{m}$ in a basis that directly resolves changes to the relative phases between the eigenstates.
@@ V Conclusions
A minimally disturbing measurement is completely characterized by the conditional probabilities $p(m|a)$ that describe the information gain with respect to the eigenstates $\left|a\right\rangle$. The unavoidable disturbance of the input state can then be identified by expanding the self-adjoint operators $\hat{M}$ into a sum of unitary operations representing the changes of the input states. As we have shown above, the maximal disturbance is described by a set of unitary operators that correspond to the Fourier transforms of the projection operators $\{\left|a\right\rangle\thinspace\left\langle a\right|\}$. This fundamental transformation of a description of information gain into a conjugated description of state transformations shows how the quantum formalism introduces a necessary relation between the information gained about one physical property and the physical changes of another. Eq. (14) describes the experimentally observable bound on information gained about $a$ for a specific pattern of disturbance observed in the complementary property $b$. This bound can be particularly useful in the detection of possible information leaks along a quantum channel. By observing the pattern of disturbance for the inputs $\{\left|b\right\rangle\}$, an upper limit can be placed on the possible leakage of information encoded in $\{\left|a\right\rangle\}$. This limit may be applied to BB84 type quantum cryptography protocols where the information is encoded in two complementary basis systems $\{\left|a\right\rangle\}$ and $\{\left|b\right\rangle\}$ [19]. Transmission errors observed in $\{\left|b\right\rangle\}$ indicate possible information leaks in the transmission of $\{\left|a\right\rangle\}$ and vice versa. Eq.(14) thus relates the transmission characteristics of the channel directly to the security of quantum cryptography protocols.
It should be emphasized that the complementary basis $\{\left|b\right\rangle\}$ appears in the analysis as a description of the distinguishable effects of unitaries $\hat{U}(k)$ with eigenstates $\left|a\right\rangle$. The relation between the basis sets $\{\left|a\right\rangle\}$ and $\{\left|b\right\rangle\}$ thus originates from the relation between phase changes and experimentally observable changes in quantum mechanics. It might be worth noting that the same fundamental relation between phases and observable output is expressed by the quantum Fourier transform that appears in Shor’s algorithm [20]. In general, the quantum formalism introduces a fundamental relation between information and dynamics by describing quantum processes as Hilbert space operators, where the same algebra is used for the representation of information gain by projection operators and for the representation of disturbance by unitary operators. The relation between information and disturbance formulated in this work thus captures a fundamental aspect of quantum mechanics with no correspondence in either classical physics or classical information theory.